OpenStudy (gabylovesyou):

@iambatman What is the completely factored form of d^4 - 8d^2 + 16? (d2 + 4)(d2 - 4) (d2 - 4)(d2 - 4) (d2 + 4)(d + 2)(d - 2) (d + 2)(d - 2)(d + 2)(d - 2)

OpenStudy (gabylovesyou):

i got d^2(d^2 - 8 + 16/d^2)

OpenStudy (acxbox22):

you can not factor out d^2 because there is no d^2 in 16

OpenStudy (jdoe0001):

hint: \(\large { d^4 - 8d^2 + 16\implies \begin{array}{cccllll} ({\color{brown}{ d^2}})^2&-8{\color{brown}{ d^2}}&+16\\ &-4-4&-4\cdot -4 \end{array} }\)

OpenStudy (welshfella):

d^4 - 8d^2 + 16 = (d^2 - 4)(d^2 - 4) now you factor d^2 - 4 - the difference of 2 squares do you recall what these fact to?

OpenStudy (welshfella):

a^2 - b^2 = (a + b)(a - b)

OpenStudy (gabylovesyou):

OHHH so its B

OpenStudy (welshfella):

no - it will factor further..

OpenStudy (acxbox22):

close but you have to factor further the (d^2-4) terms can be factored like @welshfella said

OpenStudy (gabylovesyou):

ok how do i do that then

OpenStudy (gabylovesyou):

oh wait just the symbols so A ?

OpenStudy (welshfella):

take square root of d^2 and square root of 4 - these are the a and b in the formula i posted

OpenStudy (gabylovesyou):

ok.... d and 2

OpenStudy (acxbox22):